A man pushed a 1,000 kg car to the +x direction. If the applied force is 5000N, what is the acceleration of the the coefficient of the friction force is 0.57

1
Expert's answer
2022-01-20T10:07:20-0500

Apply Newton's second law:


"ma=F_\text{net},\\
ma=F-f,\\
ma=F-\mu mg,\\\space\\
a=\dfrac{F}{m}-\mu g=-0.586\text{ m/s}^2."